OpenStudy (anonymous):

Use the quotient rule to simplify the following expression. Assume that x>0. Square root: 72x^3 / square root: 2x

OpenStudy (rc94):

so you have \[\frac{ \sqrt{72x ^{3}} }{ \sqrt{2x} }\]? Because that's the same as \[\sqrt{36x ^{2}} = \pm6x\]

OpenStudy (anonymous):

ok, awesome

OpenStudy (anonymous):

it is, but forget the \(\pm\) part, it is just plus